# Boardwalk: Towards a Framework for Creating Board Games with LLMs

### v1

@@ 1 Introduction
Board games are an important object of study in the field of Artificial Intelligence (AI), with many works concerning themselves with developing algorithms for playing or modifying games [Browne 2011, Todd et al. 2024]. However, in order for such studies to be conducted, they first require the researchers to represent the game they’re studying in code. This can be a time-consuming task, especially for games of higher complexity with various different pieces and edge cases to consider.
For this reason, several Game Description Languages (GDLs), such as Ludii [Piette et al. 2020], have been developed; these are domain-specific languages meant to facilitate the implementation and analysis of digital board games, by providing direct commands for defining the components and rules of such games. However, creating game descriptions with GDLs still requires knowledge of the language’s own syntax, which sometimes requires convoluted constructions for simple operations of specific games [^1] For example, Ultimate Tic-Tac-Toe in Ludii’s GDL has 9 levels of nested tests to check move validity [Ludii Portal 2020]..
Therefore, Large Language Models (LLMs) present themselves as highly appealing tools to aid in this task. Modern LLMs have shown great performance in code generation tasks [Coignion et al. 2024], adding significant efficiency to code-dependent workflows. One particularly interesting aspect of LLMs is their capacity for In-Context Learning (ICL), being able to perform tasks they had not been trained for simply from receiving additional information pertaining to those tasks as part of the prompt [Brown et al. 2020].
@@ 2 Related Work
Most existing studies in LLM-oriented procedural content generation for games are centered on video games. [Todd et al. 2023] were able to train LLMs to create novel and functional Sokoban levels. [Sudhakaran et al. 2023] were able to fine-tune GPT2 to create Super Mario Bros levels with custom features defined by natural language prompts. LLMs have also been used to tell interactive narratives in games [Sun et al. 2023, Kumaran et al. 2023]. All of these focused on creating content for otherwise human-implemented games.
@@
Using a video game description language, [Hu et al. 2024] had LLMs generate both rules and levels for maze-based games. They found that the models required context-rich prompts for optimal performance and faced greater challenges with the logical aspects of game design compared to the syntax of the language.
@@
[Browne 2011] reports the seemingly first automated system to generate a commercial board game from scratch [^2] Yavalath: https://boardgamegeek.com/boardgame/33767/yavalath. It uses evolutionary algorithms to evolve game rules on the Ludi (with one ‘i’) Game Description Language (GDL), combined with automated playtesting with general game playing (GGP) [Genesereth et al. 2005] algorithms. However, the Ludi framework is not available for further experimentation.
@@
The second version of Ludii (now with two ‘i’) GDL [Piette et al. 2020], had an initial focus on game archaeology and not on automated game creation. However, recent work has managed to use Ludii in automated board game creation. For example, [Todd et al. 2024] fine-tuned an LLM in the Ludii GDL. Using a fill-in-the-middle approach, the authors removed parts of pre-existing game descriptions and tasked the LLM with completing them, generating potentially different rules, without having to create new games from scratch. However, generated rules often contained introns, i.e. elements that were never used in gameplay.
@@
[Tanaka and Simo-Serra 2024] used ICL to generate Ludii implementations from natural language rule descriptions. However, due to the complexity of the Ludii grammar, they followed a process for each game of iteratively finding a minimal valid subgrammar to describe the game, and then iteratively finding a description over that grammar that correctly implements the rules. Not only was this very computationally demanding, requiring repeated inferences for each game, but it further showed that using the Ludii language requires a significant amount of upfront work to familiarize the LLM with its structure. Only 72% of descriptions generated by their method compiled (i.e. were valid in the Ludii GDL grammar).
@@
Given the results of [Todd et al. 2024, Tanaka and Simo-Serra 2024], we decided not to use a specific GDL for our purposes, due to the significant limitations in code generation that stem from the LLMs’ lack of familiarity with them. Instead, our Boardwalk GGP API (see Section 3) aims for simplicity: the games are defined as Python programs. Using a programming language is easier for LLMs as they consume extensive programming content in their pre-training [Jiang et al. 2024]. Thus, the LLMs are likely to have a significantly higher success rate writing code in a popular programming language rather than a specific GDL. Moreover, this removes an entry barrier for human programmers as well, because there is no extra learning curve for a new language.
@@ 4 Methodology
This section describes the methods used to answer our research question, i.e., whether LLMs can implement digital versions of board games from rules described in natural language. We conducted tests on three different, state-of-the-art LLMs: DeepSeekV3 [Liu et al. 2025], Claude 3.7 Sonnet [Anthropic 2024], and ChatGPT-4o [OpenAI et al. 2024]. These LLMs were accessed via the Poe website [^4] https://poe.com/. They will hereby be referred to as DeepSeek, Claude and GPT, respectively.
@@ 6 Conclusion
It is important to emphasize that these results were obtained from zero-shot inference experiments: no examples of code were provided to the models for context. In-context learning (giving several examples of the task at hand as part of the prompt) [Brown et al. 2020] as well as few-shot learning (training the LLM for subsequent steps) greatly improve LLM performance. Thus, we intend to explore these approaches in future work, as we expect them to both increase the LLMs’ understanding of the API and to better understand game descriptions in natural language.

### v2

@@ 1 Introduction
Board games are an important object of study in the field of Artificial Intelligence (AI), with many works concerning themselves with developing algorithms for playing or modifying games [Browne2011evolutionary, todd2024gavel]. However, in order for such studies to be conducted, they first require the researchers to represent the game they’re studying in code. This can be a time-consuming task, especially for games of higher complexity with various different pieces and edge cases to consider.
For this reason, several Game Description Languages (GDLs), such as Ludii [ludii], have been developed; these are domain-specific languages meant to facilitate the implementation and analysis of digital board games, by providing direct commands for defining the components and rules of such games. However, creating game descriptions with GDLs still requires knowledge of the language’s own syntax, which sometimes requires convoluted constructions for simple operations of specific games [^1] For example, Ultimate Tic-Tac-Toe in Ludii’s GDL has 9 levels of nested tests to check move validity [ultimateTictactoe]..
Therefore, Large Language Models (LLMs) present themselves as highly appealing tools to aid in this task. Modern LLMs have shown great performance in code generation tasks [Coignion2024llmcoding], adding significant efficiency to code-dependent workflows. One particularly interesting aspect of LLMs is their capacity for In-Context Learning (ICL), being able to perform tasks they had not been trained for simply from receiving additional information pertaining to those tasks as part of the prompt [few-shot-learners].
@@ 2 Related Work
Most existing studies in LLM-oriented procedural content generation for games are centered on video games. [todd2023levels] were able to train LLMs to create novel and functional Sokoban levels. [mariogpt] were able to fine-tune GPT2 to create Super Mario Bros levels with custom features defined by natural language prompts. LLMs have also been used to tell interactive narratives in games [Sun2023, Kumaran2023]. All of these focused on creating content for otherwise human-implemented games.
@@
Using a video game description language, [hu2024] had LLMs generate both rules and levels for maze-based games. They found that the models required context-rich prompts for optimal performance and faced greater challenges with the logical aspects of game design compared to the syntax of the language.
@@
[Browne2011evolutionary] reports the seemingly first automated system to generate a commercial board game from scratch [^2] Yavalath: https://boardgamegeek.com/boardgame/33767/yavalath. It uses evolutionary algorithms to evolve game rules on the Ludi (with one ‘i’) Game Description Language (GDL), combined with automated playtesting with general game playing (GGP) [Genesereth2005ggp] algorithms. However, the Ludi framework is not available for further experimentation.
@@
The second version of Ludii (now with two ‘i’) GDL [ludii], had an initial focus on game archaeology and not on automated game creation. However, recent work has managed to use Ludii in automated board game creation. For example, [todd2024gavel] fine-tuned an LLM in the Ludii GDL. Using a fill-in-the-middle approach, the authors removed parts of pre-existing game descriptions and tasked the LLM with completing them, generating potentially different rules, without having to create new games from scratch. However, generated rules often contained introns, i.e. elements that were never used in gameplay.
@@
[ludii-grammar-llm] used ICL to generate Ludii implementations from natural language rule descriptions. However, due to the complexity of the Ludii grammar, they followed a process for each game of iteratively finding a minimal valid subgrammar to describe the game, and then iteratively finding a description over that grammar that correctly implements the rules. Not only was this very computationally demanding, requiring repeated inferences for each game, but it further showed that using the Ludii language requires a significant amount of upfront work to familiarize the LLM with its structure. Only 72% of descriptions generated by their method compiled (i.e. were valid in the Ludii GDL grammar).
@@
Given the results of [todd2024gavel, ludii-grammar-llm], we decided not to use a specific GDL for our purposes, due to the significant limitations in code generation that stem from the LLMs’ lack of familiarity with them. Instead, our Boardwalk GGP API (see Section 3) aims for simplicity: the games are defined as Python programs. Using a programming language is easier for LLMs as they consume extensive programming content in their pre-training [Jiang2024llmcode]. Thus, the LLMs are likely to have a significantly higher success rate writing code in a popular programming language rather than a specific GDL. Moreover, this removes an entry barrier for human programmers as well, because there is no extra learning curve for a new language.
@@ 4 Methodology
This section describes the methods used to answer our research question, i.e., whether LLMs can implement digital versions of board games from rules described in natural language. We conducted tests on three different, state-of-the-art LLMs: DeepSeekV3 [deepseekai2025deepseekv3technicalreport], Claude 3.7 Sonnet [claude3], and ChatGPT-4o [openai2024gpt4technicalreport]. These LLMs were accessed via the Poe website [^4] https://poe.com/. They will hereby be referred to as DeepSeek, Claude and GPT, respectively.
@@ 6 Conclusion
It is important to emphasize that these results were obtained from zero-shot inference experiments: no examples of code were provided to the models for context. In-context learning (giving several examples of the task at hand as part of the prompt) [few-shot-learners] as well as few-shot learning (training the LLM for subsequent steps) greatly improve LLM performance. Thus, we intend to explore these approaches in future work, as we expect them to both increase the LLMs’ understanding of the API and to better understand game descriptions in natural language.